If you draw a card with a value of two or less from a standard deck of cards, I will pay you $463  If not, you pay me $36? (Aces are considered the highest card in the deck.)

Step 2 of 2 :  

If you played this game 712 times how much would you expect to win or lose? Round your answer to two decimal places. Losses must be entered as negative.
Answer #1

Since aces are highest card,

P(2 or less) = 4/52 = 1/13

So P(More than 2) = 1 - 1/13 = 12/13

Hence,

Expected value = $463*(1/13) - $36*(12/13) = 31/13

Therefore,

Expected winning = 712*(31/13) = 1697.85